OpenStudy (anonymous):

A rectangular prism has a base with side lengths x & x+3 and a height 2x. What is the surface area & volume of the prism? Select one: a. 5x^{2}+9x ; 2x^{3}+6x^{2} b. 10x^{2}+18 ; 2x^{3} + 6x^{2} c. 10x^{2}+18x ; 2x^{2} + 6x d. 10x^{2}+18x ; 2x^{3} + 6x^{2}

OpenStudy (bradely):

surface area,SA =2(wl+hl+hw) Here, l =x+3,w=x,and h=2x Volume,V =whl Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

So what would the answer be? im confused @bradely